a student is asked to solve b^2+x^2=a^2 and gives the solution:x=a-b. What did they do wrong and how do you correctly solve this????

Answer by edjones(8007)

b^2+x^2=a^2
x^2=a^2-b^2
x=sqrt(a^2-b^2), x=-sqrt(a^2-b^2)
.
Ed